# Batch Estimation of a Steady, Uniform, Flow-Field from Ground Velocity and Heading Measurements

### v1

@@ Abstract
This work has been submitted to IFAC for possible publication
@@ 1 Introduction
Autonomy and control algorithm can leverage estimates of the flow-field to make informed decisions that improve marine vehicle endurance, navigation, and performance. There are numerous examples in the path planning literature that demonstrate how knowledge of the flow-field can be used to plan or adjust vehicle paths for efficiency and robustness. This paper is motivated by the authors’ prior work (Wolek et al., 2021) that investigated time-optimal path planning for an underwater vehicle to re-inspect points of interest along circular orbits with a sonar sensor. In such applications, the presence of a flow-field impacts the vehicle motion and the pointing angle of a directional sonar sensor (due to sideslip). The methods presented here can be used to estimate flow direction and magnitude on-the-fly (e.g., after executing each circular inspection orbit) to support path re-planning.
@@
For a vehicle operating in a steady, uniform, flow-field the flow-relative velocity ${\bm{v}}_{\rm rel}$, the flow velocity ${\bm{w}}$, and the inertial (ground) velocity ${\bm{v}}_{\rm g}$ are related by the vector equation ${\bm{v}}_{\rm g}={\bm{v}}_{\rm rel}+{\bm{w}}$, illustrated in Fig. 1. When two of the three quantities are known, the remaining one can be determined by direct computation. For marine vehicles, ground velocity, ${\bm{v}}_{g}$, is usually provided by the navigation system by processing GPS position measurements, acoustic ranging measurements, and/or from a Doppler velocity log (DVL) sensor operating in bottom-lock mode. The vehicle flow-relative velocity ${\bm{v}}_{\rm rel}$ can also be measured with appropriate instrumentation (e.g., pitot tubes or a DVL operating in water current profiling mode). The flow velocity ${\bm{w}}$ is also often inferred as part of an inertial navigation system using various filtering techniques based on the vehicle’s kinematics (Rhudy et al., 2015) or a dynamic model (Hegrenaes and Hallingstad, 2011). Non-uniform flow-fields can be estimated, for example, using motion tomography (Chang et al., 2017).
@@
Figure 1: Left: The velocity triangle. The locus of points along the dashed circle represents the set of attainable ground velocities for a fixed magnitude $||{\bm{v}}_{\rm rel}||$. The center of the circle is defined by the flow velocity ${\bm{w}}$. Right: At least three unique ground velocity vectors are required to determine the dashed-line circle.
@@
The contributions of this paper are three batch estimation techniques to estimate water current direction and magnitude for a marine craft using noisy measurements of kinematic variables and basic model of the vehicle’s motion. The methods to do not require knowledge of the vehicle’s flow-relative speed (since it is estimated as part of the procedure) and they are relatively simple to implement with minimal tuning required. The first method is based on quadratic curve fitting to speed-over-ground magnitude and heading angle data when the vehicle executes a 360 degree heading change maneuver (e.g., a circular orbit). The second method relies on a constrained nonlinear least-square optimization to match speed-over ground velocity components and heading data. The third method is similar to the second but uses speed-over ground magnitude, rather than velocity components, for optimization. The approaches are compared to each other, as well as to another existing method in the literature based on circular curve fitting, using both simulated and experimental data.
@@
The remainder of the paper is organized as follows. Section 2 describes two curve-fitting based methods. Section 3 describes two optimization-based methods. Section 4 provides a comparison with simulated data, and Sec. 5 compares the methods using experimental data. Section 6 concludes the paper.
@@ 2 Curve-Fitting-Based Methods
Consider the following kinematic equations describing the motion of a marine craft in the horizonal plane:
@@
where $(x,y)$ is the planar position, $v$ is the speed of the vehicle, $w$ is the magnitude of the current, and $\theta$ is the current direction. This section describes two methods for estimation the constants $(v,w,\theta)$ in (1)–(2) based on curve fitting to noisy data of either ground velocity components or ground velocity magnitude as a function of heading.
@@ 2.1 Circular Curve Fit with $(\dot{x},\dot{y})$ Data
In (McLaren, 2008) it was shown that given a minimum of three unique ground velocity component measurements, $(\dot{x},\dot{y})$ from (1)–(2), the parameters $(v,w,\theta)$ can be estimated. Here, we summarize the approach from (McLaren, 2008) and use it for comparisons later on in Secs. 4 and 5.
Consider the sketch in Fig. 1 (right panel) showing the velocity triangle and a triplet of velocity triangles for three different flow-relative velocities. A circle that passes through the three ground velocity vectors can be inscribed and its center corresponds to the current velocity $(w_{x},w_{y})=(w\cos\theta,w\sin\theta)$. The radius of the circle is the vehicle’s flow-relative speed $v$. One can confirm, by substituting (1)–(2), that all points on the circle satisfy
@@
Suppose that $N$ measurements $\{\dot{x}_{i},\dot{y}_{i}\}_{i=1}^{N}$ are recorded while the vehicle performs a maneuver. In the absence of noise, the data satisfies a system of $N$ equation: ${\bm{A}}{\bm{c}}={\bm{b}}$, where
@@
and ${\bm{c}}=[-2w_{x},-2w_{y},-v^{2}+w_{x}^{2}+w_{y}^{2}]^{\rm T}=[c_{1},c_{2},c_{3}]^{\rm T}$. If, instead, the data $\{\dot{x}_{i},\dot{y}_{i}\}_{i=1}^{N}$ is corrupted by zero-mean additive Gaussian random variables with variance $\sigma^{2}$, then the system of equations can only be satisfied in a least-square sense (i.e., minimizing $J=||{\bm{A}}{\bm{c}}-{\bm{b}}||^{2}$) and ${\bm{c}}$ is estimated as ${\bm{c}}={\bm{A}}^{\rm\dagger}{\bm{b}}$, where ${\bm{A}}^{\dagger}=({\bm{A}}^{\rm T}{\bm{A}})^{-1}$ is the matrix pseudo-inverse. The flow-field parameters are then:
@@
2.2 Quadratic Curve Fit with $(v_{g},\psi)$ Data
Now, suppose that the vehicle has access to the ground speed magnitude (rather than the velocity components) along with heading angle information. That is, the data is $\{v_{g,i},\psi_{i}\}_{i=1}^{N}$ where the ground speed magnitude is
$v_{g}=f_{v_{g}}(\psi;v,w,\theta)$ $=\sqrt{\dot{x}(t)^{2}+\dot{y}(t)^{2}}$ (15)
@@ 2.2 Quadratic Curve Fit with $(v_{g},\psi)$ Data
If the heading undergoes a full revolution, then the curve $v_{g}(\psi)$ contains a maximum and a minimum (see Fig. 3). These minima and maxima can be approximated with a quadratic curve fit to estimate $(\hat{w},\hat{v},\hat{\theta})$ as described next.
Figure 3: Example of a quadratic curve fit (red lines) to $(v_{g},\psi)$ data (corresponding to the data in Fig. 2). The noise-free curve $v_{g}(\psi)$ is shown as a black line. Using the quadratic curve-fit method the estimated values are $(\hat{v},\hat{w},\hat{\theta})=$ (2.98 m/s, 0.994 m/s., 92.6 deg).
@@
$\frac{d}{d\psi}v_{g}$ $=\frac{vw\sin(\theta-\psi)}{\sqrt{v^{2}+w^{2}+2vw\cos(\theta-\psi)}}=0\;,$ (17)
which is satisfied at $\psi=\theta$ and $\psi=\theta+\pi$. The speed over ground is a maximum when the vehicle heading is aligned with the current, $v_{g}(\psi=\theta)=v+w$, and a minimum when they are opposed, $v_{g}(\psi=\theta+\pi)=v-w$. Now consider a Taylor series approximation at these two points. At the maximum:
$v_{g}$ $\approx(v+w)-\frac{vw}{\sqrt{v^{2}+2vw+w^{2}}}(\psi-\theta)^{2}$ (18)
@@
$v_{g}$ $\approx(v-w)+\frac{vw}{\sqrt{v^{2}-2vw+w^{2}}}(\psi-(\theta+\pi))^{2}$ (20)
@@
where $\delta^{-}=1/\sqrt{v^{2}-2vw+w^{2}}$. Since $\delta^{+}<\delta^{-}$ the curvature at the minimum is greater. Let $\psi_{0}^{+}$ denote a guess for the heading angle of the maximum and consider applying a second-order polynomial least-square fit to the the data on the interval $[\psi_{0}^{+}-\lambda,\psi_{0}^{+}+\lambda]$ to give the coefficients $(a^{+},b^{+},c^{+})$. Equating these coefficients to the Taylor series approximation one can solve for an estimate $(\hat{v},\hat{w},\hat{\theta})$. A similar approach can be utilized for the minimum $(a^{-},b^{-},c^{-})$ to give another estimate. To utilize more of the available data, both quadratic curve fits can be utilized as follows. Let the coordinate of the maximum and minimum be
@@ 3 Optimization-Based Methods
This section proposes two optimization-based approaches to estimate the vehicle speed and current magnitude and direction. The first method applies to $(\dot{x},\dot{y},\psi)$ data and the second method applies to $(v_{g},\psi)$ data.
@@ 3.1 Least-square Optimization with $(\dot{x},\dot{y},\psi)$ Data
$\nabla J=\left[\begin{array}[]{ccc}\frac{\partial J}{\partial v},&\frac{\partial J}{\partial w},&\frac{\partial J}{\partial\theta}\end{array}\right]^{\rm T}\;,$ (31)
@@
where $v_{\rm min}$ and $v_{\rm max}$ are user-supplied lower and upper bounds on the vehicle’s flow-relative speed. The constraints (36) also encode for the assumption $w\leq v$ that is required for the problem to be well posed. Local minima can occur along the boundaries of the constraint set since the topology of $\theta$ is not considered in (36) and local minima may also be problematic when the dataset becomes more ambiguous due to noise, or when the maneuver contains a smaller range of heading angles. To improve robustness the optimization is run from multiple start points and the lowest cost solution is selected. The initial starting points are selected as the corners of the polytope represented by the constraints and as the centroid of the polytope. An example of the cost function and iterative optimization scheme is shown in Fig. 4. The level sets of the cost function are indicative of the uncertainty in the parameter space around the estimated point.
@@
3.2 Least-square Optimization with $(v_{g},\psi)$ Data
@@
$\frac{\partial J}{\partial v}$ $=-\sum_{i=1}^{N}\frac{\left(2v+2w(c_{\psi\theta})_{i}\right)\left((v_{g})_{i}-\sqrt{\alpha_{i}}\right)}{\sqrt{\alpha_{i}}}$ (38)
$\frac{\partial J}{\partial w}$ $=-\sum_{i=1}^{N}\frac{\left(2w+2v(c_{\psi\theta})_{i}\right)\left((v_{g})_{i}-\sqrt{\alpha_{i}}\right)}{\sqrt{\alpha_{i}}}$ (39)
$\frac{\partial J}{\partial\theta}$ $=-\sum_{i=1}^{N}\frac{2vw(s_{\psi\theta})_{i}\left((v_{g})_{i}-\sqrt{\alpha_{i}}\right)}{\sqrt{\alpha_{i}}}\;,$ (40)
@@
$H_{vv}$ $=\sum_{i=1}^{N}\frac{2\left({\alpha_{i}}^{3/2}-(v_{g})_{i}w^{2}+(v_{g})_{i}w^{2}{(c_{\psi\theta})_{i}}^{2}\right)}{{\alpha_{i}}^{3/2}}$
@@
$\quad-\frac{2(c_{\psi\theta})_{i}\left((v_{g})_{i}-\sqrt{\alpha_{i}}\right)}{\sqrt{\alpha_{i}}}$
$\quad+\frac{\left(2v+2w(c_{\psi\theta})_{i}\right)\left(2w+2v(c_{\psi\theta})_{i}\right)\left((v_{g})_{i}-\sqrt{\alpha_{i}}\right)}{2{\alpha_{i}}^{3/2}}$
$H_{v\theta}$ $=\sum_{i=1}^{N}-\frac{2w(s_{\psi\theta})_{i}\left((v_{g})_{i}w^{2}-{\alpha_{i}}^{3/2}+v(v_{g})_{i}w(c_{\psi\theta})_{i}\right)}{{\alpha_{i}}^{3/2}}$
$H_{ww}$ $=\sum_{i=1}^{N}-\frac{2v(s_{\psi\theta})_{i}\left(v^{2}(v_{g})_{i}-{\alpha_{i}}^{3/2}+v(v_{g})_{i}w(c_{\psi\theta})_{i}\right)}{{\alpha_{i}}^{3/2}}$
$H_{w\theta}$ $=\sum_{i=1}^{N}-\frac{2v(s_{\psi\theta})_{i}\left(v^{2}(v_{g})_{i}-{\alpha_{i}}^{3/2}+v(v_{g})_{i}w(c_{\psi\theta})_{i}\right)}{{\alpha_{i}}^{3/2}}$
$H_{\theta\theta}$ $=\sum_{i=1}^{N}-\frac{2w(s_{\psi\theta})_{i}\left((v_{g})_{i}w^{2}-{\alpha_{i}}^{3/2}+v(v_{g})_{i}w(c_{\psi\theta})_{i}\right)}{{\alpha_{i}}^{3/2}}$
where $c_{\psi\theta}=\cos\left(\psi_{i}-\theta\right)$ and $s_{\psi\theta}=\sin\left(\psi_{i}-\theta\right)$ and with $H_{wv}=H_{vw}$, $H_{\theta v}=H_{v\theta}$, and $H_{\theta w}=H_{w\theta}$. Both the gradient and the Hessian are well-defined only if $\alpha_{i}\neq 0$. In practice, case $\alpha_{i}=0$ occur rarely and during implementation we check for and remove any such instances from the dataset.
Unlike the previous optimization case, the cost function, gradient, and Hessian depend non-linearly on the data. The summations involving $\{\dot{x}_{i},\dot{y}_{i},\psi_{i}\}_{i=1}^{N}$ must be recomputed at each iteration. As described earlier, the analytical forms of $\Delta J$ and $H$ are provided to fmincon along with constraints (36). An example of the cost function and iterative optimization sequence is shown in Fig. 5.

### v2

@@ Abstract
© 2024 Artur Wolek and James McMahon. This work has been accepted to IFAC for publication under a Creative Commons Licence CC-BY-NC-ND
@@ 1 Introduction
Autonomy and control algorithms can leverage estimates of the flow-field to improve the operation of marine vehicles, for example, in path planning to achieve greater efficiency and robustness. This paper is motivated by the authors’ prior work (Wolek et al., 2021) that investigated time-optimal path planning for an underwater vehicle to re-inspect points of interest along circular orbits with a sonar sensor. In such applications, the presence of a flow-field impacts the vehicle motion and the pointing angle of a directional sonar sensor (due to sideslip). The methods presented here can be used to estimate flow direction and magnitude on-the-fly (e.g., after executing each circular inspection orbit) to support path re-planning.
@@
For a vehicle operating in a steady, uniform, flow-field the flow-relative velocity ${\bm{v}}_{\rm rel}$, the flow velocity ${\bm{w}}$, and the inertial (ground) velocity ${\bm{v}}_{\rm g}$ are related by the equation ${\bm{v}}_{\rm g}={\bm{v}}_{\rm rel}+{\bm{w}}$, as illustrated in Fig. 1. When two of the three quantities are known, the remaining one can be determined by direct computation. For marine vehicles, ground velocity, ${\bm{v}}_{\rm g}$, is provided by the navigation system by processing GPS position measurements, acoustic ranging measurements and/or a Doppler velocity log (DVL) sensor operating in bottom-lock mode. The vehicle flow-relative velocity ${\bm{v}}_{\rm rel}$ can also be measured with appropriate instrumentation (e.g., pitot tubes or a DVL operating in water current profiling mode). The flow velocity ${\bm{w}}$ may be inferred using various filtering techniques based on the vehicle’s kinematics (Rhudy et al., 2015) or a dynamic model (Hegrenaes and Hallingstad, 2011). Non-uniform flow-fields can be estimated, for example, using motion tomography (Chang et al., 2017).
@@
Figure 1: Left: The velocity triangle. The dashed circle represents the set of attainable ground velocities for a fixed magnitude $||{\bm{v}}_{\rm rel}||$. The center of the circle is defined by the flow velocity ${\bm{w}}$. Right: At least three unique ground velocity vectors are required to determine the dashed-line circle.
@@
The contributions of this paper are three batch estimation techniques to estimate water current direction and magnitude for a marine craft using noisy measurements of kinematic variables and a basic model of the vehicle’s motion. The methods do not require knowledge of the vehicle’s flow-relative speed (since it is estimated as part of the procedure) and they are relatively simple to implement with minimal tuning required. The first method is based on quadratic curve fitting to speed-over-ground magnitude and heading angle data when the vehicle executes a 360 degree heading change maneuver (e.g., a circular orbit). The second method relies on a constrained nonlinear least-square optimization to match speed-over-ground velocity components and heading data. The third method is similar to the second but uses speed-over-ground magnitude, rather than velocity components, for optimization. The approaches are compared to each other, as well as to another existing method in the literature based on circular curve fitting, using both simulated and experimental data.
@@
The remainder of the paper is organized as follows. Section 2 describes two curve-fitting based methods. Section 3 describes two optimization-based methods. Sections 4 and  5 provide a comparison with simulated and experimental data, respectively. Section 6 concludes the paper.
@@ 2 Curve-Fitting-Based Methods
Consider the following kinematic equations describing the motion of a marine craft in the horizontal plane:
@@
where $(x,y)$ is the planar position, $v=||{\bm{v}}_{\rm rel}||$ is the flow-relative speed, $w=||{\bm{w}}||$ is the magnitude of the current, and $\theta$ is the current direction. The speed $v$ is assumed constant (e.g., corresponding to a fixed propeller speed). The current magnitude and direction, $(w,\theta)$, is assumed to be steady and uniform over the space in which a heading change maneuver is performed. This section describes two methods for estimating $(v,w,\theta)$ in (1)–(2) based on curve fitting to noisy data of either ground velocity components or ground velocity magnitude as a function of heading.
@@ 2.1 Circular Curve Fit with $(\dot{x},\dot{y})$ Data
In (McLaren, 2008) it was shown that given a minimum of three unique ground velocity component measurements, ${\bm{v}}=[\dot{x},\dot{y}]^{\rm T}$ from (1)–(2), the parameters $(v,w,\theta)$ can be estimated. Here, we summarize this approach and use it for comparisons later on in Secs. 4 and 5.
Consider the sketch in Fig. 1 (right panel) showing the velocity triangle and a triplet of velocity triangles for three different flow-relative velocities. A circle that passes through the three ground velocity vectors can be inscribed and its center corresponds to the current velocity ${\bm{w}}=[w_{x},w_{y}]^{\rm T}=[w\cos\theta,w\sin\theta]^{\rm T}$. The radius of the circle is the vehicle’s flow-relative speed $v$. One can confirm, by substituting (1)–(2), that all points on the circle satisfy
@@
Suppose that $N$ measurements $\{\dot{x}_{i},\dot{y}_{i}\}_{i=1}^{N}$ are recorded while the vehicle performs a maneuver. In the absence of noise, the data satisfies ${\bm{A}}{\bm{c}}={\bm{b}}$, where
@@
and ${\bm{c}}=[-2w_{x},-2w_{y},-v^{2}+w_{x}^{2}+w_{y}^{2}]^{\rm T}=[c_{1},c_{2},c_{3}]^{\rm T}$. If, instead, the data $\{\dot{x}_{i},\dot{y}_{i}\}_{i=1}^{N}$ is corrupted by zero-mean additive Gaussian random variables with variance $\sigma^{2}$, then the system of equations can only be satisfied in a least-square sense (i.e., minimizing $J=||{\bm{A}}{\bm{c}}-{\bm{b}}||^{2}$) where ${\bm{c}}$ is estimated as ${\bm{c}}={\bm{A}}^{\rm\dagger}{\bm{b}}$ and ${\bm{A}}^{\dagger}=({\bm{A}}^{\rm T}{\bm{A}})^{-1}$ is the matrix pseudo-inverse. The flow-field parameters are then:
@@
2.2 Quadratic Curve Fit with $(v_{\rm g},\psi)$ Data
Now, suppose that the vehicle has access to the ground speed magnitude $v_{\rm g}:=||{\bm{v}}_{\rm g}||$ (rather than the velocity components) along with heading angle information. That is, the data is $\{v_{{\rm g},i},\psi_{i}\}_{i=1}^{N}$ where
$v_{\rm g}=f_{v_{\rm g}}(\psi;v,w,\theta)$ $=\sqrt{\dot{x}(t)^{2}+\dot{y}(t)^{2}}$ (15)
@@ 2.2 Quadratic Curve Fit with $(v_{\rm g},\psi)$ Data
Again, the data is assumed to be corrupted by additive zero-mean Gaussian noise. In general this assumption is made throughout this work, however, it is worth mentioning that for a vessel continuously rotating in one direction that is not aft-fore symmetric there may be a bias present based on the direction of rotation. To mitigate this bias, the dataset could include maneuvers performed in both clockwise and counter-clockwise orientations. If the heading undergoes a full revolution, then the curve $v_{\rm g}(\psi)$ contains a maximum and a minimum (see Fig. 3). These minima and maxima can be approximated with a quadratic curve fit to estimate $(\hat{w},\hat{v},\hat{\theta})$ as described next.
Figure 3: Example of a quadratic curve fit (red lines) to $(v_{\rm g},\psi)$ data (corresponding to the data in Fig. 2). The noise-free curve $v_{\rm g}(\psi)$ is shown as a black line. Using the quadratic curve-fit method the estimated values are $(\hat{v},\hat{w},\hat{\theta})=$ (2.98 m/s, 0.994 m/s., 92.6 deg).
@@
$\frac{{\rm d}}{{\rm d}\psi}v_{\rm g}$ $=\frac{vw\sin(\theta-\psi)}{\sqrt{v^{2}+w^{2}+2vw\cos(\theta-\psi)}}=0\;,$ (17)
which is satisfied at $\psi=\theta$ and $\psi=\theta+\pi$. The speed over ground is a maximum when the vehicle heading is aligned with the current, $v_{\rm g}(\psi=\theta)=v+w$, and a minimum when they are opposed, $v_{\rm g}(\psi=\theta+\pi)=v-w$. Now consider a Taylor series approximation at these two points. At the maximum:
$v_{\rm g}$ $\approx(v+w)-\frac{vw}{\sqrt{v^{2}+2vw+w^{2}}}(\psi-\theta)^{2}$ (18)
@@
$v_{\rm g}$ $\approx(v-w)+\frac{vw}{\sqrt{v^{2}-2vw+w^{2}}}(\psi-(\theta+\pi))^{2}$ (20)
@@
where $\delta^{-}=1/\sqrt{v^{2}-2vw+w^{2}}$. Since $\delta^{+}<\delta^{-}$ the curvature at the minimum is greater. Let $\psi_{0}^{+}$ denote a guess for the heading angle of the maximum. Applying a second-order polynomial least-square fit to the the data on the interval $[\psi_{0}^{+}-\lambda,\psi_{0}^{+}+\lambda]$ to give the coefficients $(a^{+},b^{+},c^{+})$. Equating these coefficients to the Taylor series approximation one can solve for an estimate $(\hat{v},\hat{w},\hat{\theta})$. A similar approach at the minimum gives the coefficients $(a^{-},b^{-},c^{-})$ and another estimate. To utilize more of the available data, both sets of parameters can be used. Let the coordinate of the maximum and minimum be
@@ 3 Optimization-Based Methods
This section proposes two optimization-based approaches to estimate the vehicle speed and current magnitude and direction. The first method applies to $(\dot{x},\dot{y},\psi)$ data and the second method applies to $(v_{\rm g},\psi)$ data.
@@ 3.1 Least-square Optimization with $(\dot{x},\dot{y},\psi)$ Data
$\nabla J=\left[\begin{array}[]{ccc}{\partial J}/{\partial v},&{\partial J}/{\partial w},&{\partial J}/{\partial\theta}\end{array}\right]^{\rm T}\;,$ (31)
@@
where $v_{\rm min}$ and $v_{\rm max}$ are user-supplied lower and upper bounds on the vehicle’s flow-relative speed. The constraints (36) also encode for the assumption $w\leq v$ that is required for the problem to be well posed. Local minima can occur along the boundaries of the constraint set since the topology of $\theta$ is not considered in (36). Additionally, local minima may occur with noisier datasets or when the maneuver contains a smaller range of heading angles. To improve robustness the optimization is run from multiple start points and the lowest cost solution is selected. The initial starting points are selected as the corners of the polytope represented by the constraints and as the centroid of the polytope. An example of the cost function and optimization scheme is shown in Fig. 4. The level sets of the cost function are indicative of the uncertainty in the parameter space around the estimated point.
@@
3.2 Least-square Optimization with $(v_{\rm g},\psi)$ Data
@@
$\frac{\partial J}{\partial v}$ $=-\sum_{i=1}^{N}\frac{\left(2v+2w(c_{\psi\theta})_{i}\right)\left((v_{\rm g})_{i}-\sqrt{\alpha_{i}}\right)}{\sqrt{\alpha_{i}}}$ (38)
$\frac{\partial J}{\partial w}$ $=-\sum_{i=1}^{N}\frac{\left(2w+2v(c_{\psi\theta})_{i}\right)\left((v_{\rm g})_{i}-\sqrt{\alpha_{i}}\right)}{\sqrt{\alpha_{i}}}$ (39)
$\frac{\partial J}{\partial\theta}$ $=-\sum_{i=1}^{N}\frac{2vw(s_{\psi\theta})_{i}\left((v_{\rm g})_{i}-\sqrt{\alpha_{i}}\right)}{\sqrt{\alpha_{i}}}\;,$ (40)
@@
$H_{vv}$ $=\sum_{i=1}^{N}\frac{2\left({\alpha_{i}}^{3/2}-(v_{\rm g})_{i}w^{2}+(v_{\rm g})_{i}w^{2}{(c_{\psi\theta})_{i}}^{2}\right)}{{\alpha_{i}}^{3/2}}$
@@
$\quad-\frac{2(c_{\psi\theta})_{i}\left((v_{\rm g})_{i}-\sqrt{\alpha_{i}}\right)}{\sqrt{\alpha_{i}}}$
$\quad+\frac{\left(2v+2w(c_{\psi\theta})_{i}\right)\left(2w+2v(c_{\psi\theta})_{i}\right)\left((v_{\rm g})_{i}-\sqrt{\alpha_{i}}\right)}{2{\alpha_{i}}^{3/2}}$
$H_{v\theta}$ $=\sum_{i=1}^{N}-\frac{2w(s_{\psi\theta})_{i}\left((v_{\rm g})_{i}w^{2}-{\alpha_{i}}^{3/2}+v(v_{\rm g})_{i}w(c_{\psi\theta})_{i}\right)}{{\alpha_{i}}^{3/2}}$
$H_{ww}$ $=\sum_{i=1}^{N}-\frac{2v(s_{\psi\theta})_{i}\left(v^{2}(v_{\rm g})_{i}-{\alpha_{i}}^{3/2}+v(v_{\rm g})_{i}w(c_{\psi\theta})_{i}\right)}{{\alpha_{i}}^{3/2}}$
$H_{w\theta}$ $=\sum_{i=1}^{N}-\frac{2v(s_{\psi\theta})_{i}\left(v^{2}(v_{\rm g})_{i}-{\alpha_{i}}^{3/2}+v(v_{\rm g})_{i}w(c_{\psi\theta})_{i}\right)}{{\alpha_{i}}^{3/2}}$
$H_{\theta\theta}$ $=\sum_{i=1}^{N}-\frac{2w(s_{\psi\theta})_{i}\left((v_{\rm g})_{i}w^{2}-{\alpha_{i}}^{3/2}+v(v_{\rm g})_{i}w(c_{\psi\theta})_{i}\right)}{{\alpha_{i}}^{3/2}}$
where $c_{\psi\theta}=\cos\left(\psi_{i}-\theta\right)$ and $s_{\psi\theta}=\sin\left(\psi_{i}-\theta\right)$ and with $H_{wv}=H_{vw}$, $H_{\theta v}=H_{v\theta}$, and $H_{\theta w}=H_{w\theta}$. Both the gradient and the Hessian are well-defined only if $\alpha_{i}\neq 0$. In practice, the case $\alpha_{i}=0$ occurs rarely and during implementation we check for and remove any such instances from the dataset.
Unlike the previous optimization case, the cost function, gradient, and Hessian depend non-linearly on the data. The summations involving $\{\dot{x}_{i},\dot{y}_{i},\psi_{i}\}_{i=1}^{N}$ must be recomputed at each iteration. As described earlier, the analytical forms of $\Delta J$ and $H$ are provided to fmincon along with constraints (36). An example of the cost function and optimization sequence is shown in Fig. 5.